Oahu Kiki tracks the number of units purchased and sold throughout each accounting period but applies its inventory costing method perpetually at the time of each sale, as if it uses perpetual inventory system. Assume Oahu Kiki’s records show the following for the month of January. The company sold 270 units between January 16 and 23. Date Units Unit Cost Total Cost Beginning Inventory January 1 140 $ 75 $ 10,500 Purchase January 15 300 85 25,500 Purchase January 24 240 105 25,200 Required: Calculate the cost of ending inventory and the cost of goods sold using the FIFO and LIFO methods.
Answer #1

Calculate cost of goods sold and cost of ending inventory

FIFO LIFO
Cost of goods sold (140*75+120*85) = 20700 (270*85) = 22950
Cost of ending inventory 61200-20700 = 40500 61200-22950 = 38250